OpenStudy (anonymous):

The hypotenuse of a right triangle is 20 centimeters, and the shorter leg is 12 centimeters. Find the length of the other leg.

OpenStudy (anonymous):

do you know the Pythagorean Theorem?

OpenStudy (anonymous):

The hypotenuse squared is the sum of the squares of the other two sides. C^2 = A^2 + B^2 is commonly used. Can you work it out from that?

OpenStudy (anonymous):

u already know the your hyp which same as c and leg too

OpenStudy (anonymous):

\[[a2+b2=hyp] or [a2−hyp2=b2] or [2−hyp2=a2]\]